If the diameter of a circle is P and P^2/4 =2whag is the area is the circle!

Answer by greenestamps(13200)


Area = (pi)(r^2)  formula for area of a circle
     = (pi)((P/2)^2)  the radius is half the diameter, which is P
     = (pi)(P^2/4)  simplify
     = (pi)(2)  it is given that P^2/4 = 2
     = 2pi